Sixx AM Manufacturing has a target debt—equity ratio of 0.55. Its cost of equity is 19 percent, and its cost of debt is 8 percent. If the tax rate is 38 percent, what is the company's WACC? Enter your answer as a percent rounded to 2 decimal places, e.g., enter 32.16% as 32.16, not 0.3216. ______%
Answer #1

Weighted average cost of capital = Weight of debt * cost of debt * ( 1 - Tax rate) + Weight of equity * cost of debt

Debt - Equity ratio = 0.55

Debt Equity ratio = Debt / Equity

0.55 = Debt / 1

Debt = 0.55. Equity = 1

Weight of debt = 0.55 / ( 0.55 +1)

= 0.3548387

Weight of equity = 1 / ( 1 + 0.55)

= 0.64516129

So, WACC = 0.3548387 * 8% * ( 1 - 38%) + 0.64516129 * 19% *

WACC = 14.02% [ Rounded to two decimals]